The outstanding capital stock of Crane Corporation consists of 1,800 shares of $100 par value, 5% preferred, and 5,000 shares of $50 par value common.

Assuming that the company has retained earnings of $80,500, all of which is to be paid out in dividends, and that preferred dividends were not paid during the 2 years preceding the current year, state how much each class of stock should receive under each of the following conditions.

(a) The preferred stock is noncumulative and nonparticipating. (Round answers to 0 decimal places, e.g. $38,487.)

Preferred

Common

$

$


(b) The preferred stock is cumulative and nonparticipating. (Round answers to 0 decimal places, e.g. $38,487.)

Preferred

Common

$

$


(c) The preferred stock is cumulative and participating. (Round the rate of participation to 4 decimal places, e.g.1.4278%. Round answers to 0 decimal places, e.g. $38,487.)

Preferred

Common

$

$
Answer #1

Preferred Dividend = 1800 shares * $100 *5% = $9,000

a) Preferred stock is non cumulative and non-participating

Since preferred dividend is non cumulative, no arrears in dividend will be paid

So, Preferred stock dividend = $9,000

Common stock dividend ($80,500 - $9,000) = $71,500.

b) The preferred stock is Cumulative and Non Participating

Preferred dividend = ($9000*2) + $9,000 = $27,000

Common stock dividend (80,500 - 27,000) = $53,500

C) The preferred stock is Cumulative and Participating

Preferred dividend = ($9000*2) + $9,000 = $27,000

Common stock dividend ($5,000 * 50 * 5%) = $12,500

Balance dividend = $80,500 - $27,000 - $12,500 = $41,000

Participating payout

Preferred = $41,000 * (180,000 / 430,000) = $17,163

Common stock = $41,000 * (250,000 / 430,000) = $23,837

Total Preferred dividend = $27,000 + $17,163 = $44,163

Total Common stock dividend = $12,500 + $23,837 = $36,337